1.What is the wavelength of the yellow sodium emission, which has a frequency of 5.09 × 1014/s.

2.What is the frequency of violet light with a wavelength of 408 nm

3.The red spectral line of lithium occurs at 671 nm (6.71× 1014 m). Calculate the energy of one photon of this light.




1
Expert's answer
2021-03-05T03:29:30-0500

The wavelength of yellow light is 580nm = 580×10−9m.
B) The frequency of yellow light, 
ν=λ
c​=580×10−9m
3.0×108m/s​=5.17×1014/s.
The wave number of yellow light, 
C) ν-=λ1​=580×10−9m
1​=1.72×106/m.